OpenStudy (anonymous):

What is the modulus of 7 - 8i?

OpenStudy (anonymous):

How do you do this?

OpenStudy (anonymous):

what is modulus?

OpenStudy (anonymous):

hard to explain, at least for me it is

mathslover (mathslover):

@mridrik Do you know how to find the modulus of the general form of complex numbers?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

I know like 8 % 12 = 8 and stuff like that

mathslover (mathslover):

OK no problem, let me give you an introduction to the modulus of complex numbers.

mathslover (mathslover):

Let there be Z (a complex number ) = a + ib So , we have to find the modulus of Z = |Z| = |a+ib| It is given by |Z| = \(\sqrt{x^2+y^2}\)

mathslover (mathslover):

oops sorry |Z| = \(\sqrt{a^2+b^2}\)

OpenStudy (unklerhaukus):